# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ Abstract.
Cryptocurrencies and central bank digital currencies (CBDCs) are reshaping the monetary landscape, offering transparency and efficiency while raising critical concerns about user privacy and regulatory compliance. This survey provides a comprehensive and technically grounded overview of privacy-preserving digital currencies, covering both cryptocurrencies and CBDCs. We propose a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to underlying cryptographic primitives, protocol mechanisms, and system architectures. Unlike previous surveys, our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving currencies through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges at the intersection of cryptography, distributed systems, and policy definition, which motivate further investigation into the primitives and design of digital currencies that balance real-world privacy and auditability needs.
@@ 1. Introduction
The rise of cryptocurrencies and the rapid development of central bank digital currencies (CBDCs) have fueled renewed interest in digital currencies. This momentum stems from the advances in decentralized technologies (i.e., blockchains and smart contracts), the introduction of decentralized finance, and the increasing concerns over data privacy and financial surveillance. While cryptocurrencies have demonstrated that permissionless, cryptographically secure value transfer is feasible, central banks are interested in modernizing monetary infrastructures with digital-native capabilities. However, these innovations raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). Public blockchains typically offer pseudo-anonymity, using addresses decoupled from real identities. However, the transparency of public blockchains enables transactions tracing, address clustering, and ultimately users de-anonymization, thus exposing financial behavior and undermining confidentiality (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). As a result, privacy-preserving digital currencies have emerged as a critical research field, attracting attention from cryptographers, system designers, economists, and policymakers.
@@
Researchers have developed a variety of privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)) and confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native cryptocurrencies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). These solutions implement various notions of privacy, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: compliance with regulatory frameworks—such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT)—requires mechanisms for auditability and selective disclosure. Therefore, recent work has explored cryptographic protocols such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments to reconcile these conflicting goals.
@@
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital currencies, spanning both decentralized cryptocurrencies and CBDCs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. We introduce a systematic taxonomy of privacy and auditability properties and map them to concrete implementations across a broad range of systems. Structured as a hitchhiker’s guide, our survey defines key design goals and illustrates how they have been implemented in the literature through cryptographic primitives, protocols, and system architectures. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital currencies along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized in digital currency systems, mapping high-level goals to technical mechanisms. We trace the evolution of privacy-preserving digital currencies across three generations. Third, we identify research challenges stemming from the design of privacy-preserving digital currencies, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Our survey aims to organize this diverse and technically rich landscape, supporting both foundational research and practical protocol design in the rapidly evolving domain of digital currencies.
@@
This paper is organized as follows. Section 2 reviews related surveys and outlines how our work complements and differs from them. Section 3 introduces the key cryptographic primitives used in digital currencies and cryptocurrencies. Section 4 presents a detailed analysis of privacy-preserving currency designs along key architectural and cryptographic dimensions. Section 5 outlines open challenges and future research directions, and Section 6 concludes the paper.
@@ 2. Related Surveys
A large body of literature has explored privacy-preserving digital currencies, motivated by both the technological challenges and the societal implications of financial surveillance. This includes surveys on PETs, systematization of cryptocurrencies, and privacy investigations for CBDCs.
Surveys on Privacy in Cryptocurrencies
Among the earliest works, Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and introduced altcoins (i.e., alternatives to Bitcoin), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized cryptocurrencies, while emphasizing the absence of a unified definition of privacy, which hinders fair comparisons between systems. More recent surveys have attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features, and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives. While technically detailed, their coverage of cryptocurrency-specific protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution or design trade-offs.
Other related surveys take a narrower or orthogonal perspective. Almashaqbeh and Solomon (Almashaqbeh and Solomon, 2022) focus primarily on privacy in blockchain computation, i.e., smart contracts, but do not consider concerns of accountability in payments. Zhang (Zhang, 2023) and Wang et al. (Wang et al., 2020) provide useful overviews of privacy in deployed cryptocurrencies, but lack systematic analysis of technical contributions. Alsalami and Zhang (Alsalami and Zhang, 2019) present a classification of PETs but do not address accountability aspects.
@@
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms or specific cryptographic solutions. Darbha and Arora (Bank of Canada)111https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective. These works offer blueprints but lack detailed cryptographic analysis. Our work bridges this gap by providing a comparative review of cryptographic primitives, protocols, and system architectures that spans both CBDCs and cryptocurrencies.
@@
Several surveys focus on specific technologies or techniques. Examples include Bitcoin (Ghesmati et al., 2022), smart contracts (Benarroch et al., 2024; Li et al., 2022), DeFi (Baum et al., 2023), mixers (Arbabi et al., 2023), accountability mechanisms (Chatzigiannis et al., 2021), and ZKPs (Sun et al., 2021; Almashaqbeh and Solomon, 2022; Liang et al., 2025). These contributions do not consider the full design space of privacy-preserving digital currencies as we do, nor do they explore how different privacy features interact—particularly in settings requiring auditability. Koerhuis et al. (Koerhuis et al., 2020) present a forensic analysis of Monero and Verge, and highlight how implementation flaws threaten privacy; a concern we also emphasize.
@@
Compared to previous works, our survey offers a comprehensive framework for analyzing privacy-preserving digital currencies. We propose a unified taxonomy that integrates both decentralized cryptocurrencies and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
@@ 3. Background
We first introduce key cryptographic primitives, including commitments, encryption schemes, and ZKPs. Then, we briefly review key cryptographic protocols used in privacy-preserving currencies, such as anonymity-enhanced signatures, multi-party computation, and one-time public keys.
@@ 3.1. Commitments and Homomorphic Commitments
Homomorphic commitments, such as Pedersen (Pedersen, 1992) or ElGamal (Elgamal, 1985), preserve the addition or multiplication of corresponding values. For example, let and be elements of a prime group such that is not feasible to compute, let be a random value over and let be the committer’s value. Perdersen commitments are defined as . The homomorphic properties of Pedersen commitments guarantee that namely, the sum of the commitments of and is equal to the commitment of the sum of the two values under the sum of the two randomness and , allowing multiple operations on the committed messages. Pedersen commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
@@ 3.2. Encryption Schemes
3.2.1. Identity-based Encryption
Identity-based encryption (IBE) schemes (Shamir, 1985) are public-key encryption schemes in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
3.2.2. Threshold Encryption
Threshold encryption allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
3.2.3. Homomorphic Encryption
Homomorphic encryption schemes (Rivest et al., 1978) allow to compute additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the operation applied to the plaintexts, i.e., , where represent an encryption function and denotes either addition or multiplication. This allows to delegate computation to a third party while maintaining confidentiality. In payment systems (e.g., (Ma et al., 2021; Bünz et al., 2020)), it enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments.
@@ 3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling one party (the prover) to convince another party (the verifier) that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK), when the verifier is convinced through a single message. ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZK proof is short and fast to verify, it is succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the -protocol (or Sigma-protocol), which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, -protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g.,(Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproof is optimized for range proof and does not require trusted setup. zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024). We usually refer to NIZK as ZK.

### v2

# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

@@ Abstract.
Crypto-assets and central bank digital currencies (CBDCs) are reshaping how value is exchanged in distributed computing environments. These systems combine cryptographic primitives, protocol design, and system architectures to provide transparency and efficiency while raising critical challenges around privacy and regulatory compliance. This survey offers a comprehensive overview of privacy-preserving digital payment systems, covering both decentralized ledgers and CBDCs. We present a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to underlying cryptographic primitives, protocols, and system architectures. Our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving digital payment systems through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges, motivating further research into architectures and solutions that balance strong privacy with real-world auditability needs.
@@ 1. Introduction
The rapid proliferation of digital payment systems has sparked renewed interest in designing distributed architectures that balance user privacy with institutional accountability. Spanning decentralized crypto-assets111The term cryptocurrency is widely used to describe decentralized systems such as Bitcoin, Ethereum, or Monero, that enable peer-to-peer asset transfer without a central authority. However, this usage differs from the terminology adopted by central banks and regulators (ECB, 2018), which reserve the term “currency” for instruments with legal tender status. To reflect this distinction and bridge the terminology gap between these communities, we adopt the more nuanced terms crypto-assets or digital payment systems throughout the paper. and centrally managed infrastructures such as central bank digital currencies (CBDCs), these systems leverage cryptographic techniques and distributed consensus to enable secure, efficient, and programmable value transfer. However, the same transparency and auditability that make these systems attractive also raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). For example, public blockchains typically offer pseudo-anonymity. This enables transactions tracing, address clustering, and ultimately users de-anonymization, thus exposing financial behavior (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). As a result, privacy-preserving digital payments have emerged as a critical research field, attracting attention from cryptographers, engineers, economists, and policymakers.
@@
Researchers have developed a variety of privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)) and confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native crypto solutions (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). They implement various notions of privacy, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: compliance with regulatory frameworks—such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT)—requires mechanisms for auditability and selective disclosure. Therefore, recent work has explored cryptographic protocols such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments to reconcile these conflicting goals.
@@
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital payment systems, covering both decentralized ledgers and centrally managed systems with privacy guarantees—including CBDCs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital payments systems along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized, mapping high-level goals to technical mechanisms. We trace the development of privacy-preserving digital payment systems, highlighting the role of cryptographic techniques across three generations. Third, we identify open research challenges stemming from the design of privacy-preserving digital payment systems, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Supplementary material further contextualizes key publication venues and technical details of specific privacy-enhancing mechanisms.
@@
This paper is organized as follows. Section 2 reviews related surveys and outlines how our work complements and differs from them. Section 3 introduces the key cryptographic primitives used in privacy-preserving digital payments, which are then analyzed along key architectural and cryptographic dimensions in Section 4. Then, Section 5 outlines open challenges and future research directions, and Section 6 concludes the paper.
@@ 2. Related Surveys
A large body of literature has explored privacy-preserving digital payment infrastructures, driven by both the technical challenges and societal concerns about financial surveillance.
Privacy in Crypto-assets
Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and introduced altcoins (i.e., alternatives to Bitcoin), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized crypto solutions, while emphasizing the absence of a unified definition of privacy. More recent surveys attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features, and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives. While technically detailed, their coverage of payment protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution and design trade-offs.
Other related surveys take a narrower or orthogonal perspective. Almashaqbeh and Solomon (Almashaqbeh and Solomon, 2022) focus primarily on privacy in blockchain computation, i.e., smart contracts; Zhang (Zhang, 2023) and Wang et al. (Wang et al., 2020) overview privacy features in deployed public ledger-based systems; Alsalami and Zhang (Alsalami and Zhang, 2019) present an analysis of PETs, not considering accountability.
@@
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms. Darbha and Arora (Bank of Canada)222https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective. These works offer blueprints but lack detailed cryptographic analysis. Conversely, our work presents a comparative review of technical design choices including primitives, protocols, and architectures.
@@
Several surveys focus on specific technologies (e.g., Bitcoin (Ghesmati et al., 2022), smart contracts (Benarroch et al., 2024; Li et al., 2022), DeFi (Baum et al., 2023), mixers (Arbabi et al., 2023), accountability mechanisms (Chatzigiannis et al., 2021), and ZKPs (Sun et al., 2021; Almashaqbeh and Solomon, 2022; Liang et al., 2025)). These contributions do not consider the full design space of privacy-preserving payment systems as we do, nor do they explore how different privacy features interact—particularly in settings requiring auditability. Koerhuis et al. (Koerhuis et al., 2020) present a forensic analysis of Monero and Verge, and highlight how implementation flaws threaten privacy.
@@
We propose a unified taxonomy of privacy-preserving payment systems that integrates both decentralized ledger-based solutions and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
@@ 3. Background
We first introduce key cryptographic primitives (Sections 3.1–3.3), and then we briefly review key cryptographic protocols used in privacy-preserving digital payment systems (Section 3.4–3.6).
@@ 3.1. Commitments and Homomorphic Commitments
Homomorphic commitments, such as Pedersen (Pedersen, 1992) or ElGamal (Elgamal, 1985), preserve the addition or multiplication of corresponding values. In particular, Pedersen commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
@@ 3.2. Encryption Schemes
- Identity-based encryption (IBE):
schemes (Shamir, 1985) are public-key encryption schemes in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
- Threshold encryption:
allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
- Homomorphic encryption:
schemes (Rivest et al., 1978) allow to compute additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the operation applied to the plaintexts. This allows to delegate computation to a third party while maintaining confidentiality. It enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments (e.g., (Ma et al., 2021; Bünz et al., 2020)).
@@ 3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling a prover to convince a verifier that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK). ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZK proof is short and fast to verify, it is succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the -protocol, which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, -protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g., (Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproof is optimized for range proof and does not require trusted setup. Zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024).